# A probabilistic, data-driven closure model for RANS simulations with aleatoric, model uncertainty

### v1

@@ 3.1 Test case: Backward Facing step
We select the backward facing step configuration in order to assess the proposed modeling framework. This is a classic benchmark problem that has been widely used for studying the performance of turbulence models as it poses significant challenges due to the complex flow features such as flow separation, reattachment and recirculation Nadge and Govardhan (2014); Pioch et al. (2023). In this setup, as illustrated in Fig. (2) a two-dimensional channel flow is abruptly expanded into a rectangular cavity with a step change in height. The flow separates at the step and forms two recirculation zones downstream, one directly after the step and the other on the upper channel wall downstream. These two recirculation zones affect the reattachment length. The flow features can be seen in more detail as depicted in the LES simulations in Fig. (4). In this setting, the Reynolds number is defined as:
$Re=\frac{uh}{\nu}$ (37)
@@
3.2 Training data generation
@@
We used the open-source CFD platform OpenFOAM Jasak et al. (2007) for the LES and baseline RANS simulations. We utilized the steady-state, incompressible solver simpleFoam for the baseline RANS simulations. This solver uses the Semi-Implicit Method for Pressure Linked Equations (SIMPLE) in order to solve both the momentum and pressure equations. For the LES simulations, we employed the pimpleFoam transient solver, which combines both the PISO (Pressure Implicit with Split Operator) and SIMPLE algorithms to solve the pressure and momentum equations. In particular, we use the WALE (Wall-Adapting Local Eddy-Viscosity) model Nicoud and Ducros (1999). This model is well-suited for capturing the turbulent structures near solid walls and is known for its accurate predictions of wall-bounded turbulent flows. We discretized both the baseline RANS and LES domains using second-order methods and all the meshes were non-uniform with mesh density increasing in the domains of interest. To ensure numerical accuracy, we ran all simulations with a CFL number below 0.3. Other pertinent details of the LES and baseline simulations are provided in Table 2.
A few snapshots of the instantaneous velocity magnitude for $Re=1100$ from the LES simulation are depicted in Fig. (3), where one can observe how the flow features of interest evolve over time. Only the time-averaged velocities/pressures were used for training which are shown in Fig. (4). The mean field reference data from LES is interpolated to the same mesh used for the Finite Element (FE)-based calibrated RANS model (model implementation details discussed in the sequel). However, not all the observations in the grid were used for training. The density of observations was higher in the regions of steeper velocity gradients (in the reattachment and recirculation regions) and lower in other regions. In the subsequent results, mean velocity/pressure observations at approximately $8\%$ of the grid points of the RANS-FE mesh were used, i.e. a rather sparse dataset (some indicative observation points are depicted in Figure 10). The influence of the number of observation points on the learning and predictive results is an interesting research direction, which we will be addressed in future investigations.
@@
Table 2: Parameters used for performing the CFD simulations for the generation of the training and test data Gresho et al. (1993)
@@ 3.3 Differentiable forward RANS solver and probabilistic learning implementation
For the discretization of the RANS equations (Equation (12)) the Finite Element (FE) method was employed and the implementation was carried out in the open source package FEniCS Alnæs et al. (2015) due to its innate adjoint solver Mitusch et al. (2019). The basic quantities and their dimensions are listed in Table 3. Further details about the differentiable solver can be found in A.
Probabilistic inference and learning tasks were performed using the probabilistic programming package Pyro Bingham et al. (2019) which is built on top of the popular machine learning library PyTorch Paszke et al. (2019). The ELBO maximization was performed using the ADAM scheme Kingma and Ba (2014). The number of Monte Carlo samples used at each iteration for the estimation of the ELBO and its gradient was $K=5$ (Algorithm 1). The gradient computation of the ELBO (Equation (35)) was enabled by overloading the autograd functionality of PyTorch to facilitate interaction between the solver’s adjoint formulation and the auto-differentiation-based neural network gradient. A relatively small learning rate of $10^{-6}$ was employed due to the Monte Carlo noise in the ELBO gradients. The neural network architecture employed for the parametric RS model was identical to the one suggested by Ling et al. (2016b) where the optimal number of hidden layers and nodes per layer was determined to be 8 and 30 respectively. The Leaky ReLU was chosen as the activation function for all layers. We noted however that the usual practice of random weight initialization was unsuitable as it led to divergence of the solution even after applying the stabilization schemes. For this reason, we used baseline RANS closure data with added noise to pre-train the neural network in order to provide a suitable initialization.
@@ 3.4 Results and Discussion
We assessed the trained model for the test-case with $Re=500$ which was not contained in the training data. In the sequel, various aspects of the probabilistic predictions obtained as described in Section (2.2.5) are compared with the reference LES and the baseline RANS predictions. Even though the same parametrized RS closure term was used in Ling et al. (2016b) (and other subsequent works branching from this), their results are not directly comparable due to the use of blending functions Kaandorp and Dwight (2020); Geneva and Zabaras (2019), which combine baseline RS values near the walls with the constant, predicted RS in the bulk and with the amount of blending being case-dependent.
Figure 5: Comparison of the predicted components of the RS tensor $\bm{\tau}_{\bm{\theta}}$ with the LES reference values and the baseline RANS ($k-\epsilon$) for the test case with $Re=500$. The three components $\bm{\tau}_{11}$, $\bm{\tau}_{12}$ and $\bm{\tau}_{22}$ are separately compared in three different blocks. In each block, top - the LES RS tensor component contour, second - the $k-\epsilon$ RS tensor component contour, third - the predicted, parametric RS tensor $\bm{\tau}_{\bm{\theta}}$ for $\bm{\theta}=\bm{\theta}_{MAP}$, fourth - the contour plot of the absolute error between the LES RS tensor and the $\bm{\tau}_{\bm{\theta}}$ tensor component, bottom - the inferred hyper-parameter $\bm{\Lambda}$ of the (reduced) discrepancy tensor $\bm{E_{\tau}}$.
@@
  • Even though no RS observations were provided during the training, the parametric $\bm{\tau}_{\bm{\theta}}$ (i.e. neural-network based) part of the RS closure is able to capture the basic features of the reference (i.e. LES) RS field. In contrast, the baseline $k-\epsilon$ model severely under/over-estimates its magnitude and misrepresents its spatial variability.
  • The bottom row of all the three blocks in Fig. (5) depicts the inferred precision $\bm{\Lambda}$ of the (reduced) discrepancy tensor $\bm{E}_{\tau}$. As this is inversely proportional to the predictive uncertainty, we note that it attains smaller values in the regions where the parametric model $\bm{\tau}_{\bm{\theta}}$ deviates the most from the LES values (e.g. for $105<x/h<$ 10). Conversely, it attains very large values (which correspond to practically zero model discrepancies) in areas where the parametric closure model is able to correctly account for the underlying phenomena (e.g. far downstream and for all three RS components). This is expected as the flow attains an almost parabolic profile in this region, which in turn translates to reduced fluctuations in the RS tensor, making it easier for the neural network to learn.
@@
  • Even though the RS field is captured with some discrepancies, the predicted mean fields agree well with the reference LES data (Figures 6, 7 and 8, discussed in the sequel). This points to the non-uniqueness of this inverse problem solution, also reported by other works Duraisamy (2021); Brenner et al. (2022).
  • There exist two recirculation zones in the backward-facing step flow setup. The primary one forms just after the step and the secondary appears above it for Reynolds numbers close to 400 and above, for the given expansion ratio Armaly et al. (1983). As it can be seen in Fig. (6), the proposed model is able to predict the appearance of the two recirculation zones in close agreement with the LES, whereas the baseline RANS model underestimates the size of the first recirculation zone and almost completely misses the second one.
@@
  • Similarly to the stream-wise velocity, predictions for the wall-normal velocity (Fig. (7)) and the pressure (Fig. (8)) are in good agreement with the reference LES values, as opposed to the baseline RANS. In the first recirculation zone, the baseline $k-\epsilon$ is completely off, while the predicted values with the credible interval covers the reference LES. Also, the pressure predictions (Fig. (8)) identify the crucial zone where the flow reattaches to the wall (around $x/h\approx 10$), which is very difficult to predict in general. At the reattachment point, there is a transition from low-pressure in the recirculation zone to higher pressure along the wall. The posterior standard deviation at this point is also higher than in the other regions, ensuring that the reference solution is enveloped.
Figure 6: Velocity contours in x-direction ($\bm{u}$) for the test case with Re = 500. top - ground truth (LES), second - baseline $k-\epsilon$, third - the posterior predictive mean ($\bm{u}\right>$), fourth - the contour plot of the absolute error between the LES ($\bm{u}_{LES}$) and the posterior predictive mean ($\bm{u}\right>$), bottom - the standard deviation of the posterior predictive ($\sigma(\bm{u})$)
@@
Figure 8: Pressure contours ($\bm{p}$) for the test case with Re = 500. top - ground truth (LES), second - baseline $k-\epsilon$, third - the posterior predictive mean ($\bm{u}\right>$), fourth - the contour plot of the absolute error between the LES ($\bm{p}_{LES}$) and the posterior predictive mean ($\bm{p}\right>$), bottom - the standard deviation of the posterior predictive ($\sigma(\bm{u})$)
@@
As previously mentioned, observations of mean velocities/pressures at approximately $8\%$ of the total number of grid points in the FE mesh were used for training. Fig. (10) highlights this by comparing the stream-wise velocity at three different sections $x/h$. The left subfigure depicts the section in the first re-circulation zone. It can be seen that despite having very few observation points near the wall, the $u$ predictions are able to capture the backward flow in the re-circulation regions. In contrast, the baseline $k-\epsilon$ completely fails to capture it. This can be attributed to the earlier reattachment of the flow in the baseline $k-\epsilon$ case (flow reattachment discussed in the sequel). The middle and the right subfigures depict sections further downstream, with the middle being in the second re-circulation zone and the right in the flow recovery zone. It is observed that with a relatively small number of observation points, the trained model’s prediction is in agreement with the LES. Furthermore the latter is enveloped by the credible interval constructed by considering ±3$\pm$ 3$\times$ 3 × (the posterior standard deviation). This credible interval is much tighter as compared to the one in the left subfigure.
@@
Accurately capturing the recirculation zones is crucial to getting a reliable estimate of the reattachment length, which is a key parameter in the study of separated flows, such as the case here. The reattachment length is defined as the distance from the step where the flow separates to the point at which it reattaches to the surface downstream of the step. Reattachment occurs where the velocity gradient off the wall is zero, or in other words, where the wall shear stress is zero. The predicted reattachment length ($x_{reattach}$) by the proposed method is compared with a) the LES data (Section (3.2)) b) the baseline RANS (Section (3.2)) c) the two-dimensional, LES simulation performed by Biswas et al. (2004) for the expansion ratio of $1.9423$, and d) the results in Geneva and Zabaras (2019), who used the same Tensor Basis Neural Networks (TBNN) Ling et al. (2016b) employed in the parametric closure term in our work as well. The results are summarized in Table 4 where it is evident that while previous works deviated significantly from the reference LES value, our probabilistic prediction is able to envelop it.
The reattachment length is heavily dependent on correctly identifying the two recirculation zones and the baseline RANS model fails to predict the secondary recirculation zone (Fig. (6)). This might have resulted in such a low reattachment length. The estimate of the reattachment length is also low in Geneva and Zabaras (2019), which could be attributed to the lack of the stochastic, model correction term.

### v2

@@ 3.1 Test case: Backward Facing step
We select the backward facing step configuration in order to assess the proposed modeling framework. This is a classic benchmark problem that has been widely used for studying the performance of turbulence models as it poses significant challenges due to the complex flow features such as flow separation, reattachment and recirculation [81, 82]. In this setup, as illustrated in Fig. (3) a two-dimensional channel flow is abruptly expanded into a rectangular cavity with a step change in height. The flow separates at the step and forms two recirculation zones downstream, one directly after the step and the other on the upper channel wall downstream. These two recirculation zones affect the reattachment length. The flow features can be seen in more detail as depicted in the LES simulations in Fig. (6). In this setting, the Reynolds number is defined as:
$Re=\frac{uh}{\nu},$ (33)
@@
3.2 Generation of training data
@@
We used the open-source CFD platform OpenFOAM [83] for the LES and baseline RANS simulations. We utilized the steady-state, incompressible solver simpleFoam for the baseline RANS simulations. This solver uses the Semi-Implicit Method for Pressure Linked Equations (SIMPLE) in order to solve both the momentum and pressure equations. For the LES simulations, we employed the pimpleFoam transient solver, which combines both the PISO (Pressure Implicit with Split Operator) and SIMPLE algorithms to solve the pressure and momentum equations. In particular, we use the WALE (Wall-Adapting Local Eddy-Viscosity) model [84]. This model is well-suited for capturing turbulent structures near solid walls and is known for its accurate predictions of wall-bounded turbulent flows. In order to overcome the computational demands of LES, a domain decomposition approach was employed. In particular, we split the domain into 16 subdomains and leverage the CPU cores in parallel [^3] Computations were carried out in a Intel Core i9-12900K CPU.. We discretized both the baseline RANS and LES domains using second-order methods and all the meshes were non-uniform with mesh density increasing in the domains of interest. To ensure numerical accuracy, we ran all simulations with a CFL number below 0.3. Other pertinent details of the LES and baseline simulations are provided in Table 2. We note that the purpose of the LES model in the present work is to serve as the reference solver that the trained RANS model would try to approximate or match. The framework proposed does not make use of the particulars of the LES solver and one could readily use actual or DNS data (or combinations thereof) for training. The LES data might deviate from physical reality or the results obtained by DNS. Readers interested in the relative accuracy of LES simulations are directed to [85, 86].
The mean field reference data from LES is interpolated to the same mesh used for the Finite Element (FE)-based calibrated RANS model (model implementation details discussed in the sequel). However, not all the observations in the grid were used for training. At each RANS grid point, an independent, random coin was flipped with probability $10\%$ of picking the grid point. This resulted in mean velocity/pressure observations at approximately $8\%$ of the grid points to be used as training data, i.e. a rather sparse dataset (the selected observation points are depicted in Fig. (4)). Naturally, the density of RANS grid was higher near the step, so the density of observations was higher in the regions of steeper velocity gradients (in the reattachment and recirculation regions) and lower in other regions. Hence, the training dataset $\mathcal{D}=\{\hat{\bm{z}}^{(i)}\}_{m=1}^{M}$ consists of the velocities/pressure at those grid points obtained from the $M=4$ LES simulations carried out with for the corresponding $Re$ numbers (in Table 2). . The influence of the number of observation points on the learning and predictive results is an interesting research direction, which we will address in future investigations. A few snapshots of the instantaneous velocity magnitude for $Re=1100$ from the LES simulation are depicted in Fig. (5) where one can observe how the flow features of interest evolve over time. Only the time-averaged velocities/pressures were used for training which are shown in Fig. (6).
Figure 4: (Random) grid points where LES velocities/pressure were used for training data. We note that the total number of LES grid points is $\mathcal{O}(10^{5})$ whereas LES simulation data at approximately $1000$ grid points were employed.
@@
Table 2: Parameters used for performing the CFD simulations for the generation of the training and test data [51]
@@ 3.3 Differentiable forward RANS solver and probabilistic learning implementation
For the discretization of the RANS equations (Equation (12)) the Finite Element (FE) method was employed and the implementation was carried out in the open source package FEniCS [87] due to its innate adjoint solver [88]. The basic quantities and their dimensions are listed in Table 3. Further details about the differentiable solver can be found in A.
Probabilistic inference and learning tasks were performed using the probabilistic programming package Pyro [89] which is built on top of the popular machine learning library PyTorch [90]. The ELBO maximization was performed using the ADAM scheme [78]. The number of Monte Carlo samples used at each iteration for the estimation of the ELBO and its gradient was $K=5$ (Algorithm 1). The gradient computation of the ELBO (Equation (31)) was enabled by overloading the autograd functionality of PyTorch to facilitate interaction between the solver’s adjoint formulation and the auto-differentiation-based neural network gradient. A relatively small learning rate of $10^{-6}$ was employed due to the Monte Carlo noise in the ELBO gradients. The neural network architecture employed for the parametric RS model was identical to the one suggested by [15] where the optimal number of hidden layers and nodes per layer was determined to be 8 and 30 respectively. The Leaky ReLU was chosen as the activation function for all layers. We noted however that the usual practice of random weight initialization was unsuitable as it led to divergence of the solution even after applying the stabilization schemes. For this reason, we used baseline RANS closure data with added noise to pre-train the neural network in order to provide a suitable initialization.
@@ 3.4 Results and Discussion
We assessed the trained model for the test-case with $Re=500$ which was not contained in the training data. In the sequel, various aspects of the probabilistic predictions obtained as described in Section (2.2.5) are compared with the reference LES and the baseline RANS predictions. Even though the same parametrized RS closure term was used in [15] (and other subsequent works branching from this), their results are not directly comparable due to the use of blending functions [17, 22], which combine baseline RS values near the walls with the constant, predicted RS in the bulk and with the amount of blending being case-dependent.
((a)) Comparison for $\bm{\tau}_{11}$.
((b)) Comparison for $\bm{\tau}_{12}$.
Figure 7: Comparison of the predicted components of the RS tensor $\bm{\tau}_{\bm{\theta}}$ with the LES reference values and the baseline RANS ($k-\epsilon$) for the test case with $Re=500$. The three components $\bm{\tau}_{11}$, $\bm{\tau}_{12}$ and $\bm{\tau}_{22}$ are separately compared in subfigures (a), (b), and (c) respectively (subfigure (c) in the next page). In each block, top - the LES RS tensor component contour, second - the $k-\epsilon$ RS tensor component contour, third - the predicted, parametric RS tensor $\bm{\tau}_{\bm{\theta}}$ for $\bm{\theta}=\bm{\theta}_{MAP}$, fourth - the contour plot of the absolute error between the LES RS tensor and the $\bm{\tau}_{\bm{\theta}}$ tensor component, bottom - the inferred hyper-parameter $\bm{\Lambda}$ of the (reduced) discrepancy tensor $\bm{E_{\tau}}$.
((a)) Comparison for $\bm{\tau}_{22}$.
Figure 8: (cont. from the previous page) Comparison of the predicted components of the RS tensor $\bm{\tau}_{\bm{\theta}}$ with the LES reference values and the baseline RANS ($k-\epsilon$) for the test case with $Re=500$. The three components $\bm{\tau}_{11}$, $\bm{\tau}_{12}$ and $\bm{\tau}_{22}$ are separately compared in subfigures (a), (b) and (c) respectively (subfigure (a) and (b) in the next page). In each block, top - the LES RS tensor component contour, second - the $k-\epsilon$ RS tensor component contour, third - the predicted, parametric RS tensor $\bm{\tau}_{\bm{\theta}}$ for $\bm{\theta}=\bm{\theta}_{MAP}$, fourth - the contour plot of the absolute error between the LES RS tensor and the $\bm{\tau}_{\bm{\theta}}$ tensor component, bottom - the inferred hyper-parameter $\bm{\Lambda}$ of the (reduced) discrepancy tensor $\bm{E_{\tau}}$.
@@
  • Even though no RS observations were provided during the training, the parametric $\bm{\tau}_{\bm{\theta}}$ (i.e. neural-network based) part of the RS closure is able to capture the basic features of the reference (i.e. LES) RS field. In contrast, the baseline $k-\epsilon$ model severely under/over-estimates its magnitude and misrepresents its spatial variability. We observe that some regions have relatively higher errors (for e.g., around $105<x/h<$ 10 in Fig. (8(a)) ), which is attributed to the parametric model’s inability to provide adequate closure. This model error is exactly what we attempt to capture with our proposed stochastic discrepancy term.
  • The bottom row of all the three subfigures in Fig. (8) depicts the inferred precision $\bm{\Lambda}$ of the (reduced) discrepancy tensor $\bm{E}_{\tau}$. As this is inversely proportional to the predictive uncertainty we note that it attains smaller values in the regions where the parametric model $\bm{\tau}_{\bm{\theta}}$ deviates the most from the LES values (e.g. for $105<x/h<$ 10). Conversely, it attains very large values (which correspond to practically zero model discrepancies) in areas where the parametric closure model is able to correctly account for the underlying phenomena (e.g. far downstream and for all three RS components). This is expected as the flow attains an almost parabolic profile in this region, which in turn translates to reduced fluctuations in the RS tensor, making it easier for the neural network to learn.
@@
  • Even though the RS field is captured with some discrepancies, the predicted mean fields agree well with the reference LES data (Figures 9, 10 and 11, discussed in the sequel). This points to the non-uniqueness of this inverse problem solution, also reported by other works [7, 37].
  • There exist two recirculation zones in the backward-facing step flow setup. The primary one forms just after the step and the secondary appears above it for Reynolds numbers close to 400 and above, for the given expansion ratio [91]. As it can be seen in Fig. (9), the proposed model is able to predict the appearance of the two recirculation zones in close agreement with the LES, whereas the baseline RANS model underestimates the size of the first recirculation zone and almost completely misses the second one.
@@
  • Similarly to the stream-wise velocity, predictions for the wall-normal velocity (Fig. (10)) and the pressure (Fig. (11)) are in good agreement with the reference LES values, as opposed to the baseline RANS. In the first recirculation zone, the baseline $k-\epsilon$ is completely off, while the predicted values with the credible interval cover the reference LES. Also, the pressure predictions (Fig. (11)) identify the crucial zone where the flow reattaches to the wall (around $x/h\approx 10$), which is very difficult to predict in general. At the reattachment point, there is a transition from low-pressure in the recirculation zone to higher pressure along the wall. The posterior standard deviation at this point is also higher than in the other regions, ensuring that the reference solution is enveloped.
Figure 9: Velocity contours in x-direction ($\bm{u}$) for the test case with Re = 500. top - ground truth (LES), second - baseline $k-\epsilon$, third - the posterior predictive mean ($\bm{u}\right>$), fourth - the contour plot of the absolute error between the LES ($\bm{u}_{LES}$) and the posterior predictive mean ($\bm{u}\right>$), bottom - the standard deviation of the posterior predictive ($\sigma(\bm{u})$).
@@
Figure 11: Pressure contours ($\bm{p}$) for the test case with Re = 500. top - ground truth (LES), second - baseline $k-\epsilon$, third - the posterior predictive mean ($\bm{u}\right>$), fourth - the contour plot of the absolute error between the LES ($\bm{p}_{LES}$) and the posterior predictive mean ($\bm{p}\right>$), bottom - the standard deviation of the posterior predictive ($\sigma(\bm{u})$).
@@
As previously mentioned, observations of mean velocities/pressures at approximately $8\%$ of the total number of grid points in the FE mesh were used for training. Fig. (13) highlights this by comparing the stream-wise velocity at three different sections $x/h$. The left subfigure depicts the section in the first re-circulation zone. It can be seen that despite having very few observation points near the wall, the $u$ predictions are able to capture the backward flow in the re-circulation regions. In contrast, the baseline $k-\epsilon$ completely fails to capture it. This can be attributed to the earlier reattachment of the flow in the baseline $k-\epsilon$ case (flow reattachment discussed in the sequel). The middle and the right subfigures depict sections further downstream, with the middle being in the second re-circulation zone and the right in the flow recovery zone. It is observed that with a relatively small number of observation points, the trained model’s prediction is in agreement with the LES. Furthermore, the latter is enveloped by the credible interval constructed by considering ±3$\pm$ 3$\times$ 3 × (the posterior standard deviation). This credible interval is much tighter as compared to the one in the left subfigure.
The primary motivation behind RANS models and of this work is to reduce the computational cost of flow simulations, especially in cases where LES or DNS are prohibitively expensive. In Table 4 we report the computational time of a single, model run as compared to the baseline RANS model(for $Re=500$). We observe that even though the proposed differentiable solver is of comparable cost to the baseline RANS, it can provide far more accurate, probabilistic predictions of the LES model’s outputs which is roughly $6000$ times more computationally demanding.
LES proposed RANS solver
Relative cost 7679 1.3
Table 4: Relative computational cost in terms of computational time (on an Intel Core i9-12900K CPU) of a single model run for LES and of the proposed, differentiable RANS solver. These are compared with the cost of the baseline RANS model for Reynolds number $500$.
@@
Accurately capturing the recirculation zones is crucial to getting a reliable estimate of the reattachment length, which is a key parameter in the study of separated flows, such as the case here. The reattachment length is defined as the distance from the step where the flow separates to the point at which it reattaches to the surface downstream of the step. Reattachment occurs where the velocity gradient off the wall is zero, or in other words, where the wall shear stress is zero. The predicted reattachment length ($x_{reattach}$) by the proposed method is compared with a)the LES data (Section (3.2)) b) the baseline RANS (Section (3.2)) c) the two-dimensional, LES simulation performed by [92] for the expansion ratio of $1.9423$, and d) the results in [22], who used the same Tensor Basis Neural Networks (TBNN) [15] employed in the parametric closure term in our work as well. The results are summarized in Table 5 where it is evident that while previous works deviated significantly from the reference LES value, our probabilistic prediction is able to envelop it. The reattachment length is heavily dependent on correctly identifying the two recirculation zones and the baseline RANS model fails to predict the secondary recirculation zone (Fig. (9)). This might have resulted in such a low reattachment length. The estimate of the reattachment length is also low in [22], which could be attributed to the lack of the stochastic, model correction term.